graph each system or equations. then determine whether the system has no solution, one solution, or infinitely many solutions. if the system has one solution, name it.
2x-y=1
y=-3

Answer by rapaljer(4671)

2x-y=1
y=-3

Substitute y=-3 into the first equation:
2x-y= 1
2x -(-3) = 1
2x+ 3 = 1
2x = -2
x=-1

Solution is the point (-1,-3).

R^2 at SCC